How do you graph the linear equation x-3y=6?
Answer by checkley71(8403)

what you need are two points and then draw a straight line through these two points.
setting y=o we get x-3*0=6 or x=6 & setting x=0 we get 0-3y=6 or -3y=6 or
y=-6/3 or y=-2
thus the two points are (6,0) & (0,-2) now draw the line